ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 20 Kondratiev, Anatolii S., Institute of Mathematics and Mechanics UB RAS Finite linear groups of small degree Well-known result of M. Aschbacher on the subgroups of the finite classical groups reduces the classical problem of describing the finite linear groups of small degree to the case of absolutely irreducible quasisimple groups. We will disscuss the classification of the conjugacy classes and the normalizers of absolutely irreducible quasisimple subgroups in GL_{n}(q) for n ≤ 27. Some applications of the classification will be also presented. Part of these results is mentioned in the paper [1]. The work is partially supported by the Russian Foundation for Fundamental Investigations, grant 96-01-00488. Reference: [1] A. S. Kondratiev, Finite linear groups of small degree, In: Proc. Conf. “The Atlas ten years on”, London Math. Soc. Lect. Note Ser., to appear.